# A logic of co-valuations

### v1

@@ 3. Constructing compact spaces from posets
Let $\mathbb{P}$ be a level-injective $\omega$-chain. Then $p\mapsto p^{\in}$ is an order-isomorphism between $\bigsqcup\mathbb{P}_{n}$ and $\bigsqcup\mathbb{P}_{n}^{\in}$ ordered by inclusion.
Let $A$, $B$ be covers of a set $X$, and let $a\in A$, $b\in B$. We define $b\lhd_{B}a$ if, for $^{\prime}\in B, bb$$^{\prime}\cap b\neq\emptyset$ implies $b\subseteq b^{\prime}$; we write $B\lhd A$, when the relation $\lhd_{B}\subseteq B\times A$ is surjective, i.e., $A$ is $\rhd_{B}$-refined by $B$.
@@
Using the Lebesgue number, can easily prove that if $X$ is second-countable compact metrizable space, then for every sequence of minimal covers $\mathbb{P}_{n}$ of $X$, $n\in\mathbb{N}$, such that every $\mathbb{P}_{n}$ consolidates $\mathbb{P}_{n+1}$, and every open cover of $X$ is refined by some $\mathbb{P}_{n}$, the $\omega$-chain $(\bigsqcup\mathbb{P}_{n},\supseteq\upharpoonright\mathbb{P}_{n}\times\mathbb{P}_{n+1})$ is regular.
@@ 4. A logic of co-valuations
In the sequel, a co-bijective relation $v\subseteq X\times Y$ is called a co-valuation on $Y$ in $X$. The family of all co-valuations on $Y$ in $X$ is denoted by ${\rm{CoV}}(X,Y)$. For $v\in{\rm{CoV}}(X,Y)$, the set $X$ will be also referred to as $X_{v}$.
@@
We denote $C_{\geq}(w)=\geq\circ w$, and call $C_{\geq}(w)$ a consolidation of $w$ with pattern $\geq$, and $w$ a fragmentation of $C_{\geq}(w)$. Note that a co-valuation $v$ on $M$ is a consolidation of a co-valuation $w$ on $M$ iff the cover $[v]$ of $M$ is a consolidation of the cover $[w]$.
A sequence $(v_{i})$, $v_{i}\in{\rm{CoV}}(X_{i},M)$, of co-valuations is called an $\omega$-chain if each $v_{i}$ consolidates $v_{i+1}$, i.e., $([v_{i}],\supseteq\upharpoonright[v_{i}]\times[v_{i+1}])$ is an $\omega$-chain as a poset.
@@
Let us select a countably infinite set of variables with a special variable *. A context is a finite set of variables. In practice, we may regard any finite set as a context. Formulas $\phi(X)$ in context $X$ are generated by the grammar
$\phi(X)::=\bot\mid R\mid\neg\phi(X)\mid\phi(X)\land\psi(X)\mid\exists^{\scriptscriptstyle{f}}_{\geq}\phi(Y),$
where $\geq\in{\rm{CoV}}(X,Y)$, and $R$ is a relation symbol of arity $a(R)$ with variables contained in $X$. We write $\top$ for $\neg\bot$, $\phi\vee\psi$ for $\neg(\neg\phi\wedge\neg\psi)$, $\phi\rightarrow\psi$ for $\psi\vee\neg\phi$, $\forall^{\scriptscriptstyle{f}}_{\geq}\phi$ for $\neg\exists^{\scriptscriptstyle{f}}_{\geq}\neg\phi$, and $\exists^{\scriptscriptstyle{f}}_{X\geq Y}\phi$ to explicitly indicate the domain and co-domain of $\geq$. Note that $\exists^{\scriptscriptstyle{f}}_{X\geq Y}\phi$ has a unique context $X$ (as opposed to, e.g., atomic formulas that can have multiple contexts). A type ($n$-type) $p(X)$ is a set of formulas in a fixed signature and context $X$ (of size $n$). A type $p(X)$ is complete if for every formula $\phi(X)$ either $\phi\in p(X)$ or $\neg\phi\in p(X)$. A sentence is a formula in the context $\bar{*}=\{*\}$, and a theory is a type consisting only of sentences.
@@
Let $X$, $M$ be sets, and let $V$ a family of co-valuations on $M$. We write $V^{X}$ for $V\cap{\rm{CoV}}(X,M)$, $\bar{x}]_{v}$ for $\prod_{k}[x_{k}]_{v}$, where $\bar{x}=(x_{k})$, $x_{k}\in X$. Also, for $v\in{\rm{CoV}}(X,M)$, $w\in{\rm{CoV}}(Y,M)$, $(x_{k})\geq(y_{k})$ if $x_{k}]_{v}\supseteq[y_{k}]_{w}$, for all relevant $k$.
@@
    a family $V$ of co-valuations on $M$ in contexts that is closed under consolidations by contexts.
@@
For proto-structures $(M,V)$, $(N,W)$ in the same signature, a bijection $f:M\to N$ is an isomorphism if it preserves relations, and valuations, i.e., for $f[v]=\{(x,f(a)):(x,a)\in v\}$, where $v\in V$, and $f[V]=\{f[v]:v\in V\}$, we have $f[V]=W$.
Given a proto-structure $M=(M,V)$, a co-valuation $v\in V^{X}$, and a formula $\phi(X)$, the satisfaction relation $M\vDash_{v}\phi$ is defined by the recursion scheme, where $\bar{x}$ is a tuple in $X$:
$M\nvDash_{v}\bot,$
@@
$M\vDash_{v}\phi\wedge\psi\iff M\vDash_{v}\phi\mbox{ and }M\vDash_{v}\psi,$
@@
The type $\rm{tp}(v)$ of a co-valuation $v\in V^{X}$ is the set of all formulas $\phi(X)$ in context $X$ such that $M\models_{v}\phi$. A type $p(X)$ is realized in $(M,V)$ if there is $v\in V^{X}$ such that $M\vDash_{v}\phi$ for every $\phi\in p$; otherwise, $p$ is omitted by $M$.
@@
Let $(M,V)$ be a proto-structure in signature $L$, and let $R\in L$. Let $v\in V^{X}$, $w\in V^{Y}$, and $\geq\in{\rm{CoV}}(X,Y)$. If $M\vDash_{w}R(\bar{y})$, $w\leq v$, and $\bar{x}\in X^{a(R)}$ is such that $\bar{y}\leq\bar{x}$, then $M\vDash_{v}R(\bar{x})$.
A structure in signature $L$ is a proto-structure $(M,\{R_{\alpha}\},V)$ in signature $L$ that satisfies the following axioms:
@@
    for all $R\in L$, $v\in V^{X}$, and $\bar{x}\in X^{a(R)}$ such that $M\vDash_{v}R(\bar{x})$, there is $w\in V^{Y}$, $\bar{y}\in Y^{a(R)}$, and $\bar{a}\in[\bar{y}]_{\pi_{w}}$ such that $\bar{x}\geq^{v}_{w}\bar{y}$, and $R^{M}(\bar{a})$.
@@
The somewhat technical Axiom (II) is needed for the proof of the downward Löwenheim-Skolem theorem (Theorem 5.3) to work. Let us mention that the seemingly more natural form of Axiom (II), mimicking the requirement that co-valuations give rise to minimal covers:
@@
would be too restrictive for our purposes. For example, suppose that $R$ is a unary relation, $R^{M}=\{a_{0}\}$ is a singleton, and $v\in V^{X}$ is such that there are two distinct $x,x^{\prime}\in X$ with $a_{0}\in[x]_{v},[x^{\prime}]_{v}$. Then $M\models_{v}R(x)$ but $a_{0}\not\in[x]_{\pi_{v}}$, i.e., there is no $a\in[x]_{\pi_{v}}$ such that $R^{M}(a)$.
@@
Let $(M,V)$ be a structure. Then $V$ has amalgamation in the following sense. Let $v_{0},v_{1}\in V$, and let $\geq_{0}$, $\geq_{1}$ be such that $C_{\geq_{0}}(v_{0})=C_{\geq_{1}}(v_{1})$. Then there exist $w\in V$, and $\succeq_{0}$, $\succeq_{1}$ such that $v_{0}=C_{\succeq_{0}}(w)$, $v_{1}=C_{\succeq_{1}}(w)$, and $\geq_{0}\circ\succeq_{0}=\geq_{1}\circ\succeq_{1}$. Moreover, the latter is true for all such $w$, and $\succeq_{0}$, $\succeq_{1}$.
@@
for every structure $(M,V)$ in signature $L$, and $v\in V^{X}$.
@@
First observe that for every structure $(M,V)$ in signature $L$, and $v\in V^{Y}$
$M\models_{C_{\geq}(v)}R(\bar{x})\iff M\models_{v}\bigvee\{R(\bar{y}):\bar{x}\geq\bar{y}\},$
so the statement of the theorem holds for atomic formulas. The rest of the proof goes by induction on the complexity of formulas. For the connectives $\neg$ and $\wedge$, the inductive step is straightforward. Fix a formula of the form $\exists^{\scriptscriptstyle{f}}_{\succeq}\phi$, where $\succeq\in{\rm{CoV}}(X,W)$. Fix a structure $(M,V)$, and $v\in V^{Y}$. If $M\vDash_{C_{\geq}(v)}\exists^{\scriptscriptstyle{f}}_{\succeq}\phi$, then, by Axiom (I), there is $w\in V^{Z}$, $\geq_{0}\in{\rm{CoV}}(Y,Z)$, and $\succeq_{0}\in{\rm{CoV}}(W,Z)$ such that $M\vDash_{C_{\succeq_{0}}(w)}\phi$, and $C_{\geq_{0}}(w)=v$. Moreover, by Proposition 4.5, we have
@@
Now, we observe that, by Remark 4.6 and Proposition 4.7, we can choose $Z$ of size that is commonly bounded for every $(M,V)$. This means that there are only finitely many ways of choosing $\geq_{0}$, $\succeq_{0}$ as above. Let $\geq_{i}$, $\succeq_{i}$, $i\leq n$, enumerate all of them, and let $\psi_{i}$ be formulas given by the inductive assumption for $\phi$, $w$, and $\succeq_{i}$, $i\leq n$. Then the formula
@@
As it is more often used than $\exists^{\scriptscriptstyle{f}}_{\geq}$, especially in topological settings, we will denote it simply by $\exists_{\geq}$. The following corollary shows that $\exists^{\scriptstyle{r}}_{\geq}$ is definable in the logic of co-valuations.
@@
Let $L$ be a signature, let $X$, $Y$ be contexts, let $\geq\in{\rm{CoV}}(X,Y)$, and let $\phi(Y)$ be an formula. There is a formula $\psi(X)$ such that
@@
A structure $(M,V)$ is called an $\omega$-structure if there exists an $\omega$-chain $(v_{i})$ of admissible co-valuations generating $V$, i.e., $V=\langle\{v_{i}\}\rangle$, i.e., every admissible co-valuation is a consolidation of some $v_{i}$. We will call such $(v_{i})$ a core of $V$. As bijections preserve consolidations, we have:
@@
If $\omega$-structures $(M,V)$, $(N,W)$ have the same signature, $(v_{i})$, $(w_{i})$ are cores of $(M,V)$, $(N,W)$, respectively, $f:M\to N$ is a bijection that preservs relations, and $f[v_{i}]=w_{i}$, $i\in\mathbb{N}$, then $f$ is an isomorphism of $(M,V)$ and $(N,W)$.

### v2

@@ 3. Constructing compact spaces from posets
Let $(\mathbb{P},\geq)$ be a level-injective $\omega$-chain. Then $p\mapsto p^{\in}$ is an order-isomorphism between $(\mathbb{P},\geq)$ and the level-injective $\omega$-chain $(\bigsqcup\mathbb{P}_{n}^{\in},\supseteq)$.
Let $A$, $B$ be covers of a set $X$, and let $a\in A$, $b\in B$. We define $b\lhd_{B}a$ if, for $^{\prime}\in B, bb$$^{\prime}\cap b\neq\emptyset$ implies $b^{\prime}\subseteq a$; we write $A\rhd B$, when the relation $\rhd_{B}\subseteq A\times B$ is surjective, i.e., $A$ is $\rhd_{B}$-refined by $B$.
@@
Using the Lebesgue number, can easily prove that if $X$ is second-countable compact metrizable space, then for every sequence of minimal covers $\mathbb{P}_{n}$ of $X$, $n\in\mathbb{N}$, such that every $\mathbb{P}_{n}$ consolidates $\mathbb{P}_{n+1}$, and every open cover of $X$ is refined by some $\mathbb{P}_{n}$, the $\omega$-chain $(\bigsqcup\mathbb{P}_{n},\supseteq)$ is regular.
@@ 4. A logic of co-valuations
A co-bijective relation $v\subseteq X\times Y$ is called a co-valuation on $Y$ in $X$. The family of all co-valuations on $Y$ in $X$ is denoted by ${\rm{CoV}}(X,Y)$. For $v\in{\rm{CoV}}(X,Y)$, the set $X$ will be also referred to as $X_{v}$.
@@
We denote $C_{\geq}(w)=\geq\circ w$, and call $C_{\geq}(w)$ a consolidation of $w$ with pattern $\geq$, and $w$ a fragmentation of $C_{\geq}(w)$ with pattern $\geq$. Also, $v$ is a consolidation of $w$ (and $w$ is a fragmentation of $v$) if there exists a pattern $\geq\in{\rm{CoV}}(X,Y)$ such that $v=C_{\geq}(w)$. Note that $v$ is a consolidation of $w$ iff the cover $[v]$ of $M$ is a consolidation of the cover $[w]$ as defined in Section 3. If $v$ is a consolidation of $w$, by Proposition 3.4, we have that $\geq^{v}_{w}\in{\rm{CoV}}(X,Y)$.
A sequence $(v_{i})$, $v_{i}\in{\rm{CoV}}(X_{i},M)$, of co-valuations is called an $\omega$-chain if each $v_{i}$ consolidates $v_{i+1}$, i.e., $([v_{i}],\supseteq\upharpoonright[v_{i}]\times[v_{i+1}])$ is an $\omega$-chain poset as in Proposition 3.5.
@@
Let us select a countably infinite set of variables with a special variable *. A context is a finite set of variables. In practice, we may regard any finite set as a context.
Formulas $\phi(X)$ in context $X$ are defined by the following recursion scheme. If $R$ is a relation symbol of arity $a(R)$ with variables contained in $X$, then
$R,\,\top$
are formulas in context $X$. If $\phi(X)$, $\psi(X)$ are formulas in context $X$, then
$\neg\phi(X),\,\phi(X)\land\psi(X)$
are formulas in context $X$. Finally, if $\phi(Y)$ is a formula in context $Y$, and $\geq\in{\rm{CoV}}(X,Y)$, then
$\exists^{\scriptscriptstyle{f}}_{\geq}\phi(Y)$
is a formula in context $X$.
We write $\bot$ for $\neg\top$, $\phi\lor\psi$ for $\neg(\neg\phi\land\neg\psi)$, $\phi\rightarrow\psi$ for $\psi\lor\neg\phi$, $\forall^{\scriptscriptstyle{f}}_{\geq}\phi$ for $\neg\exists^{\scriptscriptstyle{f}}_{\geq}\neg\phi$, and $\exists^{\scriptscriptstyle{f}}_{X\geq Y}\phi$ to explicitly indicate the domain and co-domain of $\geq$. Note that $\exists^{\scriptscriptstyle{f}}_{X\geq Y}\phi$ has a unique context $X$ (as opposed to, e.g., atomic formulas that can have multiple contexts). A type ($n$-type) $p(X)$ is a set of formulas in a fixed signature and context $X$ (of size $n$). A type $p(X)$ is complete if for every formula $\phi(X)$ either $\phi\in p(X)$ or $\neg\phi\in p(X)$. A sentence is a formula in the context $\bar{*}=\{*\}$, and a theory is a type consisting only of sentences.
@@
Let $V$ be a family of co-valuations on a set $M$. We write $V^{X}$ for $V\cap{\rm{CoV}}(X,M)$, $\bar{x}]_{v}$ for $\prod_{k}[x_{k}]_{v}$, where $\bar{x}=(x_{k})$, $x_{k}\in X$. Also, for $v\in{\rm{CoV}}(X,M)$, $w\in{\rm{CoV}}(Y,M)$, $(x_{k})\geq(y_{k})$ if $x_{k}]_{v}\supseteq[y_{k}]_{w}$, for all relevant $k$.
@@
    a family $V$ of co-valuations on $M$ in contexts that is closed under consolidations by contexts, i.e., $v\in V^{Y}$, $\geq\in{\rm{CoV}}(X,Y)$ implies $C_{\geq}(v)\in V^{X}$, where $X$, $Y$ are contexts.
@@
For a proto-structure $(M,V)$, and co-valuation $v\in V^{X}$, the type $\rm{tp}(v)$ of $v$ is the set of all formulas $\phi(X)$ in context $X$ such that $M\models_{v}\phi$. For proto-structure $(N,W)$ in the same signature as $(M,V)$, a bijection $f:M\to N$ is an isomorphism if it preserves relations, and valuations, i.e., for $f[v]=\{(x,f(a)):(x,a)\in v\}$, where $v\in V$, and $f[V]=\{f[v]:v\in V\}$, we have $f[V]=W$.
Given a proto-structure $M=(M,V)$, a co-valuation $v\in V^{X}$, and a formula $\phi(X)$, the satisfaction relation $M\vDash_{v}\phi$ is defined by the following recursion scheme, where $\bar{x}$ is a tuple in $X$:
$M\vDash_{v}\top,$
@@
$M\vDash_{v}\phi\land\psi\iff M\vDash_{v}\phi\mbox{ and }M\vDash_{v}\psi,$
@@
Example 4.3.
Assume that $(M,V)$ is a proto-structure, $X=\{x_{0},x_{1}\}$ is a context, $R(x_{0},x_{1})$ is a binary relation symbol, $S(x_{1})$ is a unary relation symbol, $v\in V^{X}$, and $a_{0},a_{1},a^{\prime}_{1}\in M$. Then $a_{0},a_{1},a^{\prime}_{1}\in M$ witness that
$M\vDash_{v}R(x_{0},x_{1})\land S(x_{1})$
if $a_{0}\in[x]_{v}$, $a_{1},a^{\prime}_{1}\in[y]_{v}$, and $R^{M}(a_{0},a_{1})$, $S^{M}(a^{\prime}_{1})$. Note that it is not required that $a_{1}=a^{\prime}_{1}$.
Example 4.4.
Assume that $(M,V)$ is a proto-structure, $X=\{x_{0},x_{1}\}$, $Y=\{y_{0},y_{1},y_{2}\}$ are contexts, $\geq\in{\rm{CoV}}(X,Y)$, and $R(y_{0},y_{1},y_{2})$ is a ternary relation symbol. Consider the formula $\exists^{f}_{\geq}R(y_{0},y_{1},y_{2})$ in context $X$, and $v\in V^{X}$. A co-valuation $w\in V^{Y}$ witnesses that
$M\vDash_{v}\exists^{f}_{\geq}R(y_{0},y_{1},y_{2})$
if $v=C_{\geq}(w)$, and
$M\vDash_{w}R(y_{0},y_{1},y_{2}),$
i.e., there exist $a_{0}\in[y_{0}]_{w}$, $a_{1}\in[y_{1}]_{w}$, $a_{2}\in[y_{2}]_{w}$ such that
$R^{M}(a_{0},a_{1},a_{2}).$
@@
Let $(M,V)$ be a proto-structure in signature $L$, and let $R\in L$. Let $v\in V^{X}$, $w\in V^{Y}$, and $\geq\in{\rm{CoV}}(X,Y)$. If $M\vDash_{w}R(\bar{y})$, and $\bar{x}\in X^{a(R)}$ is such that $\bar{y}\leq\bar{x}$, then $M\vDash_{v}R(\bar{x})$.
To develop a well-behaved model theory for the logic of co-valuations, we need to work with proto-structures that satisfy certain additional requirements. A structure in signature $L$ is a proto-structure $(M,\{R_{\alpha}\},V)$ in signature $L$ that satisfies the following axioms:
@@
    for all $R\in L$, $v\in V^{X}$, and $\bar{x}\in X^{a(R)}$ such that $M\vDash_{v}R(\bar{x})$, there are $w\in V^{Y}$, $\bar{y}\in Y^{a(R)}$, and $\bar{a}\in[\bar{y}]_{\pi_{w}}$ such that $\bar{x}\geq^{v}_{w}\bar{y}$, and $R^{M}(\bar{a})$.
@@
For a structure $(M,V)$, and co-valuation $v\in V^{X}$, the type $\rm{tp}(v)$ of $v$ is the set of all formulas $\phi(X)$ in context $X$ such that $M\models_{v}\phi$. We say that a type $p(X)$ is realized in $(M,V)$ if there is $v\in V^{X}$ such that $M\vDash_{v}\phi$ for every $\phi\in p$; otherwise, $p$ is omitted by $M$. A type is consistent if there exists a structure $(M,V)$ that realizes it. A structure that realizes a theory $T$ is called a model of $T$.
@@
The somewhat technical Axiom (II) is needed for the proof of the downward Löwenheim-Skolem theorem (Theorem 5.4) to work. Let us point out that the seemingly more natural form of Axiom (II), mimicking the requirement that co-valuations give rise to minimal covers:
@@
would be too restrictive for our purposes. For example, suppose that $R$ is a unary relation symbol, $R^{M}=\{a_{0}\}$ is a singleton, and $v\in V^{X}$ is such that there are two distinct $x,x^{\prime}\in X$ with $a_{0}\in[x]_{v},[x^{\prime}]_{v}$. Then $M\models_{v}R(x)$ but $a_{0}\not\in[x]_{\pi_{v}}$, i.e., there is no $a\in[x]_{\pi_{v}}$ such that $R^{M}(a)$.
@@
Let $(M,V)$ be a structure. Then $V$ has amalgamation in the following sense. Let $v_{0},v_{1}\in V$, and let $\geq_{0}$, $\geq_{1}$ be such that $C_{\geq_{0}}(v_{0})=C_{\geq_{1}}(v_{1})$. Then there exist $w\in V$, and $\succeq_{0}$, $\succeq_{1}$ such that $v_{0}=C_{\succeq_{0}}(w)$, $v_{1}=C_{\succeq_{1}}(w)$, and
$\geq_{0}\circ\succeq_{0}=\geq_{1}\circ\succeq_{1}.$
Moreover, the above equality holds for all such $w$, and $\succeq_{0}$, $\succeq_{1}$.
@@
for every structure $(M,V)$ in signature $L$, and $v\in V^{Y}$.
@@
First observe that, for every structure $(M,V)$ in signature $L$, $v\in V^{Y}$, and $\geq\in{\rm{CoV}}(X,Y)$,
$M\models_{C_{\geq}(v)}R(\bar{x})\iff M\models_{v}\bigvee\{R(\bar{y}):\bar{y}\in Y^{a(R)},\,\bar{x}\geq\bar{y}\},$
so the statement of the theorem holds for atomic formulas. The rest of the proof goes by induction on the complexity of formulas. For the connectives $\neg$ and $\land$, the inductive step is straightforward. Fix a formula of the form $\exists^{\scriptscriptstyle{f}}_{\succeq}\phi$, where $\succeq\in{\rm{CoV}}(X,W)$. Fix a structure $(M,V)$, and $v\in V^{Y}$. If $M\vDash_{C_{\geq}(v)}\exists^{\scriptscriptstyle{f}}_{\succeq}\phi$, then, by Axiom (I), there is a context $Z$, $w\in V^{Z}$, $\geq_{0}\in{\rm{CoV}}(Y,Z)$, and $\succeq_{0}\in{\rm{CoV}}(W,Z)$ such that $M\vDash_{C_{\succeq_{0}}(w)}\phi$, and $C_{\geq_{0}}(w)=v$. Moreover, by Proposition 4.7, we have
@@
Now, we observe that, by Remark 4.8 and Proposition 4.9, we can choose $Z$ of size that is commonly bounded for all $(M,V)$, and $v\in V^{Y}$. This means that there are only finitely many ways of choosing $\geq_{0}$, $\succeq_{0}$ as above. Let $\geq_{i}$, $\succeq_{i}$, $i\leq n$, enumerate all of them, and let $\psi_{i}$ be formulas given by the inductive assumption for $\phi$, $w$, and $\succeq_{i}$, $i\leq n$. Then the formula
@@
As it is more often used than $\exists^{\scriptscriptstyle{f}}_{\geq}$, especially in topological settings, we will denote it simply by $\exists_{\geq}$ (as well as write $\forall_{\geq}\phi$ for $\neg\exists_{\geq}\neg\phi$). The following corollary shows that $\exists^{\scriptstyle{r}}_{\geq}$ is definable in the logic of co-valuations.
@@
Let $L$ be a signature, let $X$, $Y$ be contexts, let $\geq\in{\rm{CoV}}(X,Y)$, and let $\phi(Y)$ be a formula. There is a formula $\psi(X)$ such that
@@
A structure $(M,V)$ is called an $\omega$-structure if there exists an $\omega$-chain $(v_{i})$ of admissible co-valuations generating $V$, i.e., $V=\langle\{v_{i}\}\rangle$, i.e., every admissible co-valuation is a consolidation of some $v_{i}$ by a context. We will call such $(v_{i})$ a core of $V$. As bijections preserve consolidations, we have:
@@
If $\omega$-structures $(M,V)$, $(N,W)$ have the same signature, $(v_{i})$, $(w_{i})$ are cores of $(M,V)$, $(N,W)$, respectively, $f:M\to N$ is a bijection that preserves relations, and $f[v_{i}]=w_{i}$, $i\in\mathbb{N}$, then $f$ is an isomorphism of $(M,V)$ and $(N,W)$.